OpenStudy (anonymous):

factorise. (x + y)^2 + 8(x + y)

OpenStudy (anonymous):

\[(x + y)^2 + 8(x + y)= (x+y) (x+y +8) \]

OpenStudy (anonymous):

Practice factoring problems on http://www.saab.org/mathdrills/factor.cgi

OpenStudy (skullpatrol):

@Danndann if you had $$\Huge a^2+8a$$ how would you factor it?

OpenStudy (usukidoll):

just yank out the (x+y)

OpenStudy (anonymous):

@skullpatrol a(a + 8) ?

OpenStudy (skullpatrol):

Correct. Now let a = x+ y

OpenStudy (skullpatrol):

$$\Huge (x + y)^2 + 8(x + y)$$ $$\Huge \text{Let } a= x+y.$$ $$\Huge \text{Then } (x + y)^2 + 8(x + y) = a^2+8a.$$